Find the difference of two squares. The problem is (3a-b)^2-(2a-b)^2.....Thank in advance.
Answer by nyc_function(2741)

(3a-b)^2-(2a-b)^2 =
(3a - b)(3a - b) - (2a - b)(2a - b) =
9a^2 - 6ab + b^2 - (4a^2 - 4ab + b^2) =
9a^2 - 6ab + b^2 - 4a^2 + 4ab - b^2 =
5a^2 - 2ab